### v2

# Causal2Vec: Improving Decoder-only LLMs as Versatile Embedding Models

Abstract
footnotetext: Equal Contribution
Decoder-only large language models (LLMs) are increasingly used to build embedding models that effectively encode the semantic information of natural language texts into dense vector representations for various embedding tasks. However, many existing methods primarily focus on removing the causal attention mask in LLMs to enable bidirectional attention, potentially undermining the model’s ability to extract semantic information acquired during pretraining. Additionally, leading unidirectional approaches often rely on extra input text to overcome the inherent limitations of causal attention, inevitably increasing computational costs. In this work, we propose Causal2Vec, a general-purpose embedding model tailored to enhance the performance of decoder-only LLMs without altering their original architectures or introducing significant computational overhead. Specifically, we first employ a lightweight BERT-style model to pre-encode the input text into a single Contextual token, which is then prepended to the LLM’s input sequence, allowing each token to capture contextualized information even without attending to future tokens. Furthermore, to mitigate the recency bias introduced by last-token pooling and help LLMs better leverage the semantic information encoded in the Contextual token, we concatenate the last hidden states of Contextual and EOS tokens as the final text embedding. In practice, Causal2Vec achieves state-of-the-art performance on the Massive Text Embeddings Benchmark (MTEB) among models trained solely on publicly available retrieval datasets, while reducing the required sequence length by up to 85% and inference time by up to 82% compared to best-performing methods.
1 Introduction
Text embedding models encode natural language text into dense vector representations that capture contextual semantic information (text1, ; text2, ), enabling a wide range of downstream natural language processing (NLP) tasks such as information retrieval, semantic textual similarity, and question answering (karpukhin2020dense, ; xiong2021approximate, ; mteb, ). Moreover, embedding-based retrievers play a crucial role in enhancing the capabilities of large language model (LLM)-based Retrieval-Augmented Generation (RAG) systems (rag, ; liu2024chatqa, ).
For many years, pretrained language models based on encoder-only or encoder-decoder Transformer architectures, such as BERT (bert, ), RoBERTa (roberta, ), and T5 (t5, ), have been the dominant paradigm for building text embedding models. With recent advances in decoder-only LLMs, considerable efforts have focused on adapting these decoder-only architectures for embedding tasks. However, the use of causal attention in Transformer decoders leads to incomplete information encoding for each token except the last one, significantly limiting the model’s representational capacity. To address this issue, many existing LLM-based text embedding models (llm2vec, ; grit, ; nvembed, ) achieve bidirectional attention by removing the causal attention mask, enabling each token to access the entire sequence and thereby generating rich contextualized representations. Despite notable progress, we argue that modifying the original attention mechanisms of LLMs leads to a pre-train/fine-tune attention mismatch, which may compromise the model’s ability to extract semantic information acquired during pretraining. Moreover, since attention implementations vary across different LLMs (llama, ; minimax, ), replacing them may require a deep understanding of the underlying codebase and substantial modifications, which could hinder the community’s adoption of LLMs for text embedding and limit their feasibility and reliability in real-world applications. Additionally, most leading causal attention-based methods echo ; bgeicl compensate for missing contextual information in the original sequence by introducing additional textual input, which inevitably increases computational costs during training and inference, restricting their applicability in resource-constrained scenarios.
To derive text embeddings from the output hidden states of LLMs, there are two mainstream strategies: mean pooling and last-token pooling. Mean pooling involves averaging all token representations in the output sequence, whereas last-token pooling typically uses the last hidden state of the EOS token. Prior studies (zhang2023language, ; bgeicl, ) show that for approaches based on causal attention, (weighted) mean pooling is less effective than last-token pooling, since autoregressive modeling prevents earlier tokens from attending to future tokens, leading to biased aggregated representations. Consequently, last-token pooling is more commonly used in unidirectional models. However, recent works (echo, ; nvembed, ) reveal that last-token pooling can be highly sensitive to noisy information due to its reliance on tokens near the end of sequence, hindering the model’s ability to learn robust representations.
Based on the above discussions, we propose Causal2Vec, a simple yet powerful causal attention-based embedding model that significantly enhances the text encoding capabilities of decoder-only LLMs, while circumventing the need to modify their original architectures or introduce significant computational overhead. Specifically, to address the representational bottleneck inherent in causal attention mechanism while preserving the LLMs’ ability to extract semantic information learned during pretraining, we first employ a lightweight, off-the-shelf bidirectional encoder to distill the contextual content of the input text into a single Contextual token, which is then aligned to the dimensionality of LLM’s word embedding space via a trainable MLP layer. By prepending this token to LLM’s input sequence, we enable each token to access contextualized information even under the constraints of causal attention mask, without switching to bidirectional attention or utilizing extra input text. Moreover, we concatenate the last hidden states of Contextual and EOS tokens as the final text embedding, effectively mitigating the recency bias introduced by last-token pooling and encouraging LLMs to better leverage the contextualized information encoded in the Contextual token. We conduct comprehensive experiments on the MTEB benchmark (mteb, ) by integrating Causal2Vec into three decoder-only LLMs with parameter sizes ranging from 1.3B to 7B (S-LLaMA-1.3B, LLaMA-2-7B, Mistral-7B). Evaluation across 56 datasets spanning 7 tasks demonstrates that our best model, Causal2Vec-Mistral-7B, achieves state-of-the-art (SOTA) embedding performance among models trained solely on publicly available retrieval data. Notably, compared to leading causal attention-based methods echo ; bgeicl , our model reduces the required sequence length by up to 85% and inference time by up to 82%. Furthermore, we empirically conduct extensive analysis to validate the effectiveness and necessity of the proposed mechanism. Overall, our empirical results demonstrate the effectiveness of decoder-only LLMs in producing high-quality contextualized text embeddings, underscoring the significant potential inherent in the models’ original causal attention mechanism and contributing to future advancements in this research area.
2 Related Work
2.1 Bidirectional Text Embedding Models
Text embeddings have been widely applied in a variety of downstream applications such as fact checking (fact, ), similarity search (search, ), open-domain question answering (karpukhin2020dense, ), and retrieval-augmented generation (RAG) (rag, ). Over the past few years, embedding methods based on pretrained language models with bidirectional attention, such as BERT (bert, ), RoBERTa (roberta, ), and T5 (t5, ), have dominated text embedding tasks. Early notable approaches including SimCSE (simcse, ) and Sentence-T5 (st5, ), are pretrained with a masked language modeling objective and finetuned in a contrastive manner with natural language inference (NLI) datasets. Later works like E5 (e5, ) and GTE (gte, ) further improve embedding performance through weakly supervised contrastive training on curated text pair datasets. More recent methods (instrut1, ; instrut2, ; instrut3, ) have shifted toward developing general-purpose embedding models through task instructions, demonstrating strong generalization to unseen tasks.
2.2 Decoder-only LLM-Based Text Embedding Models
Since LLMs represent the most advanced language models available and have demonstrated excellent performance across a wide range of language tasks, recent research has increasingly focused on developing embedding models based on decoder-only architectures. Luo et al. luo-etal-2024-large highlight that larger models with extensive pretraining consistently improve embedding performance, showing the effectiveness of leveraging LLMs as backbone models for embedding tasks. An intuitive approach to converting decoder-only LLMs into text encoders is to utilize the last hidden state of the EOS token under causal attention as text embedding. Ma et al.repllama finetuned LLaMA-2 to serve as both a dense retriever and a point-wise reranker, demonstrating that LLMs can indeed outperform smaller models in retrieval tasks. Llama2Vecllama2vec is an unsupervised learning method that enhances dense retrieval by reconstructing the input sentence and leveraging text embeddings to predict the next sentence. Despite these advances, decoder-only LLM-based embedding methods still suffer from inherent architectural drawbacks: causal attention prevents each token from interacting with subsequent tokens, hindering the model’s ability to produce contextual representations. To address this limitation, LLM2Vec (llm2vec, ) transformed the LLM’s attention from unidirectional to bidirectional by removing the causal attention mask. Building upon this approach, NV-Embed (nvembed, ) introduced a novel latent attention layer over the final hidden states, followed by mean pooling to generate higher-quality representations. Moreover, GRITLM (grit, ) unifies embedding tasks and generative tasks via different attention mechanisms.
Notably, the aforementioned bidirectional LLM-based embedding methods involve modifications to the model architecture, which may not be compatible with various LLM backbones, thereby significantly limiting their generality in real-world applications. In contrast, some studies preserve the original causal attention mechanism while attempting to address the inherent limitation of decoder-only LLMs. SGPT (sgpt, ) proposed a position-weighted mean pooling strategy that assigns higher weights to tokens at later positions, serving as an alternative to last-token pooling or regular mean pooling. E5-Mistral (e5mistral, ) finetuned decoder-only LLMs using synthetic data and task-specific instructions. ECHO (echo, ) repeated the input twice in the autoregressive modeling paradigm, allowing the text embedding extracted from the repeated tokens to capture contextualized information. Additionally, PromptEOL (prompteol, ) and bge-en-icl (bgeicl, ) enhance text embeddings by leveraging the in-context learning (ICL) capabilities of LLMs, augmenting the original input with task-specific examples to provide contextual information. Similarly, our work focuses on improving the text embedding quality of decoder-only LLMs by increasing the information density of each token within the sequence, thereby enabling the final EOS token to capture richer semantic information about the entire text. However, there are several key differences: 1) We employ a lightweight bidirectional encoder to generate the Contextual token without introducing additional overhead to LLMs, whereas ECHO and bge-en-icl suffer from significantly increased computational cost at both training and inference time. 2) We concatenate the last hidden states of Contextual and EOS tokens as the final text embedding, effectively mitigating the recency bias caused by last-token pooling and empowering LLMs to better understand the contextual information encoded in the Contextual token. As a result, our approach achieves state-of-the-art performance on the MTEB benchmark among models that train only on publicly available retrieval data, showing its simplicity and effectiveness in text embedding tasks.
3 Method
Figure 1 illustrates the overall pipeline of our proposed Causal2Vec in generating text representations. Given an input text, we first utilize the lightweight bidirectional encoder to produce a Contextual token. This token is concatenated with the word embeddings of a task-specific instruction, the input text, and the EOS token to form the final input sequence, which is then fed into the LLM for causal attention computation. The final text embedding is obtained by concatenating the output hidden states of Contextual and EOS tokens. We elaborate on the Contextual token and representation method in the following sections.
Figure 1: Overview of our proposed Causal2Vec method.
3.1 The Contextual Token
The remarkable capacity of large language models (LLMs) for human language understanding and generation is acquired through training on massive amounts of text data (ouyang2022training, ; wei2022chain, ; brown2020language, ), showcasing their effectiveness in encoding semantic information. However, the inherent causal attention in LLMs prevents earlier tokens from accessing information about future tokens, thus limiting the model’s representational capacity. As a result, many LLM-based embedding models switch from unidirectional attention to bidirectional by removing the causal attention mask (llm2vec, ; grit, ; nvembed, ). Although bidirectional attention facilitates effective information flow across the entire sentence, it introduces an attention mismatch between pretraining and finetuning, which somewhat compromises the LLMs’ ability to extract semantic information acquired during pretraining.
To fully unlock the potential of LLMs for text embedding, it is essential to address the limitations of causal attention while preserving LLMs’ ability to extract well-learned semantic information. To this end, we first introduce a lightweight BERT-style model that encodes the input text into a dense vector representation $h\in\mathbb{R}^{1\times k}$, termed the "Contextual Token". Specifically, this token is generated by applying mean pooling over the last hidden states of the additional bidirectional model, capturing contextualized information about the entire input. Furthermore, to bridge the gap between the BERT-style model and LLM, we employ a simple MLP layer to match the dimensionality of the Contextual token with LLM’s word embedding space, and then encourage LLM to understand the sentence information encoded in this token through contrastive learning. Motivated by llava2 , the MLP layer consists of two linear transformations with a GELU activation $\sigma$, which can be formulated as:
$C=\sigma(h\boldsymbol{\mathrm{W}}_{1}^{\top})\boldsymbol{\mathrm{W}}_{2}^{\top},$ (1)
where $\boldsymbol{\mathrm{W}}_{1}\in\mathbb{R}^{d\times k}$ and $\boldsymbol{\mathrm{W}}_{2}\in\mathbb{R}^{d\times d}$ are trainable projection matrices, and $C\in\mathbb{R}^{1\times d}$ denotes the language embedding of the Contextual token, which shares the same dimensionality $d$ as LLM’s word embedding space. Moreover, to leverage LLMs’ instruction-follow capabilities for producing general-purpose text embeddings, we use task-specific instructions for both training and evaluation, following  (e5mistral, ; llm2vec, ; echo, ). Consequently, by adding the instruction, Contextual and EOS tokens to the start and end of input sequence, the resulting sequence fed into LLM can be constructed as:
$\boldsymbol{\mathrm{x}}=[\boldsymbol{\mathrm{I}};C;\boldsymbol{\mathrm{T}};EOS]\in\mathbb{R}^{(l+n+2)\times d},$ (2)
where $\cdot$$\cdot$] denotes the vertical concatenation operation, $\boldsymbol{\mathrm{I}}\in\mathbb{R}^{l\times d}$ and $\boldsymbol{\mathrm{T}}\in\mathbb{R}^{n\times d}$ represent the word embeddings of the task-specific instruction and input text, respectively. In this way, each token following the Contextual token $C$ can capture contextualized information even without attending to future tokens. More importantly, the use of Contextual token requires no modifications to the model architecture, which not only preserves LLMs’ ability to extract semantic information learned during pretraining, but also enables seamless integration across different LLMs.
3.2 Representation Method
Last-token pooling typically utilizes the final hidden state of the EOS token as text embedding, and has been widely adopted in unidirectional models (zhang2023language, ; e5mistral, ; bgeicl, ), since only the last EOS token captures information from the entire input. However, recent studies (echo, ; nvembed, ) indicate that the EOS token embedding depends heavily on the hidden states of tokens near the end of sequence, leading to potential semantic bias in long-text scenarios.
To address this issue, we introduce a simple yet effective representation method tailored for the proposed embedding framework. Specifically, we concatenate the last hidden states of Contextual and EOS tokens as the final sentence representation for general-purpose embedding generation. Recognizing that the Contextual token has initially captured the semantic content of the input text, concatenating it with the EOS token yields a vector representation with richer contextualized content. Moreover, this approach enables explicit supervision of the Contextual token during training, thereby helping LLMs better understand the semantic information encoded in this added token.
The proposed representation method is optimized through supervised contrastive learning with the standard InfoNCE loss (loss, ), which can be formulated as:
$\mathcal{L}=-\log\frac{\exp(f(q,p^{+})/\tau)}{\exp(f(q,p^{+})/\tau)+\sum_{j=1}^{N}\exp(f(q,p_{j}^{-})/\tau)},$ (3)
where $f(q,p^{+})$ represents the scoring function that computes cosine similarity between the query-passage pair embeddings from retrieval datasets, $p^{-}$ denotes both in-batch and hard negative examples, and $\tau$ is a temperature hyperparameter fixed at 0.05 in all experiments.
4 Experiments
4.1 Training Data
For training, we follow the mainstream practices (e5mistral, ; llm2vec, ; nvembed, ; bgeicl, ), utilizing a collection of publicly available retrieval datasets curated by echo , which consists of approximately 1.5M samples. More details about the training dataset composition can be found in Appendix A.3.
Notably, many previous works achieve strong performance by using extensive in-domain non-retrieval datasets from MTEB, private datasets, or proprietary synthetic data. To ensure fairness and consistency in comparisons, we only evaluate models trained on public retrieval datasets, enabling the verification of models’ generalization capability to unseen non-retrieval tasks, which serves as a critical criterion for defining a general-purpose embedding model.
4.2 Evaluation
For evaluation, we use the English subset of the Massive Text Embeddings Benchmark (MTEB) (mteb, ), which comprises 56 datasets spanning 7 embedding task categories, aiming to verify whether the model can produce universal text embeddings suitable for various embedding tasks. Specifically, the task categories include Retrieval (Retr.), Reranking (Rerank.), Clustering (Clust.), Pair Classification (PairClass.), Classification (Class.), Semantic Textual Similarity (STS), and Summarization (Summ.). The main evaluation metrics are nDCG@10, MAP, V-measure (V-meas.), average precision (AP), accuracy (Acc.), and Spearman correlation (Spear., both for STS and Summ.), respectively.
The full evaluation of a 7B parameter model on the MTEB benchmark is GPU resource intensive, requiring over 300 A100 80GB GPU hours. To speed up the evaluation, we follow llm2vec and echo . Specifically, we introduce the MTEB-MINI for ablation studies and analysis, which combines 41 datasets covering all task categories in MTEB. Details of the MTEB-MINI composition can be found in Appendix B.1.
4.3 Implementation Details
For the base model, we integrate the proposed Causal2Vec with three decoder-only LLMs with parameter sizes ranging from 1.3B to 7B, including Sheared-LLaMA-1.3B (S-LLaMA-1.3B) (sheared, ), Llama-2-7B-chat (LLaMA-2-7B) (llama2, ), and Mistral-7B-Instruct-v0.2 (Mistral-7B) (mistral, ). Regrading the off-the-shelf bidirectional encoder, we adopt E5-base-v2 (e5, ), a lightweight model with only 110M parameters. All LLMs are finetuned using low-rank adaptation (LoRA) (hu2022lora, ) on A100 80GB GPUs. In particular, LoRA is also applied to the bidirectional encoder when using the 1.3B LLM (refer to section 4.5.2 for the reason). Refer to Appendix A and B for further experimental details.
Table 1: Performance comparison on the full MTEB for models trained on publicly available retrieval data. Scores are averaged across each task category. S-LLaMA-1.3B, LLaMA-2-7B, and Mistral-7B refer to embedding methods built upon these decoder-only LLMs. The models grouped under Miscellaneous show reported scores from various recent works that use an encoder and/or a decoder other than the aforementioned LLMs. The best results are highlighted in bold, and the second-best are underlined. See Appendix C.5 for detailed results for each dataset.
Task (# of datasets) Retr. (15) Rerank. (4) Clust. (11) PairClass. (3) Class. (12) STS (10) Summ. (1) Avg (56)
Metric nDCG@10 MAP V-Meas. AP Acc. Spear. Spear.
Miscellaneous
[…]
S-LLaMA-1.3B
LLM2Vec llm2vec 51.44 55.38 43.57 86.20 72.21 83.58 30.01 61.85
ECHO echo - - - - - - - 62.01
Causal2Vec 52.69 56.54 44.35 86.18 72.94 83.76 31.45 62.63
LLaMA-2-7B
LLM2Vec llm2vec 54.60 57.38 45.24 88.03 76.33 83.73 28.49 64.14
Causal2Vec 55.28 58.18 47.23 87.85 75.95 84.90 31.09 64.94
Mistral-7B
[…]
ECHO echo 55.52 58.14 46.32 87.34 77.43 82.56 30.73 64.68
GRITLM grit 53.10 61.30 48.90 86.90 77.00 82.80 29.40 64.70
LLM2Vec llm2vec 55.99 58.42 45.54 87.99 76.63 84.09 29.96 64.80
NV-Embed nvembed 59.00 59.59 45.44 87.59 73.93 79.07 30.16 64.18
MGH mgh 57.49 58.80 47.96 87.83 77.62 84.04 31.10 65.87
bge-en-icl (zero-shot) bgeicl 59.59 56.85 42.61 87.87 75.47 83.30 29.52 64.67
bge-en-icl (few-shot) bgeicl 60.08 56.67 46.55 88.51 77.31 83.69 30.68 66.08
Causal2Vec 57.28 59.46 48.89 88.43 76.41 85.38 30.57 66.10
Causal2Vec (w/ ICL) 57.48 59.36 50.78 89.19 77.53 85.66 30.82 66.85
4.4 MTEB Results
We evaluate the proposed Causal2Vec against recent state-of-the-art text embedding methods on the full MTEB benchmark. It is worth noting that many existing approaches rely on non-public or synthetic data for training. Even among methods limited to disclosed data, many use substantial amounts of non-retrieval MTEB-related data, potentially leading to overfitting  (bgeicl, ). To ensure fair comparisons and better verify the model’s generalizability across diverse embedding tasks, we compare only against models trained on publicly available retrieval data. As shown in Table 1, Causal2Vec demonstrates consistently strong performance across various LLM backbones, with our best model, Causal2Vec-Mistral-7B, achieving state-of-the-art results on par with bge-en-icl (bgeicl, ). Surprisingly, although LLaMA-2-7B has been shown to perform significantly worse than Mistral-7B on embedding tasks grit ; llm2vec , our Causal2Vec built upon LLaMA-2-7B surpasses most leading Mistral-7B-based methods, except for bge-en-icl (few-shot).
Comparison to Bidirectional LLM-based Methods. LLM2Vec (llm2vec, ) removes the causal attention mask to enable bidirectional attention, demonstrating competitive performance on the MTEB benchmark. NV-Embed (nvembed, ) builds upon this approach by introducing a latent attention layer to obtain pooled embeddings, addressing the limitations of mean pooling. Notably, NV-Embed achieves significant improvements by incorporating additional non-retrieval MTEB-related and synthetic data. However, when trained solely on public retrieval data, NV-Embed shows lower performance compared to LLM2Vec. GRITLM (grit, ) unifies embedding and generation training paradigms, yielding embedding performance on par with LLM2Vec. In contrast to these bidirectional LLM-based methods, our Causal2Vec requires no modifications to the model architecture, yet enables each token in the sequence to access contextual information through the introduced Contextual token. More importantly, our method consistently outperforms LLM2Vec across various base models under identical training data, achieving improvements of $0.78$ points for S-LLaMA-1.3B, $0.80$ for LLaMA-2-7B, and $1.30$ for Mistral-7B. These results underscore that shifting from causal attention to bidirectional attention is not necessary for adopting LLMs to text embedding tasks—and may even compromise the model’s ability to extract the well-learned semantic information. We argue that the effectiveness of bidirectional attention in capturing contextual information relies on maintaining consistency in attention mechanism throughout both pretraining and finetuning.
Comparison to ECHO. ECHO echo repeats the input, allowing each token from the second occurrence to access the complete sequence content. However, this strategy comes with a significant drawback: it doubles the maximum sequence length, thus increasing computational cost during both training and inference. Conversely, our Causal2Vec enables contextual information access through a single Contextual token, without introducing additional overhead to the LLM. Furthermore, our proposed representation method, which concatenates the the last hidden states of Contextual and EOS tokens as final text embedding, effectively mitigates the recency bias inherent in last-token pooling used by ECHO and reduces reliance on the EOS token. As a result, Causal2Vec consistently outperforms ECHO on S-LLaMA-1.3B and Mistral-7B by $0.62$ points and $1.42$ points, confirming the effectiveness of our design in leveraging LLMs’ causal attention for text embedding.
Comparison to bge-en-icl. bge-en-icl (bgeicl, ) endows LLM-based embedding models with in-context learning (ICL) (icl, ) capabilities by incorporating multiple task-related examples into the input. Our method significantly outperforms bge-en-icl (few-shot) on clustering, reranking, and STS tasks with an improvement of at least $1.69$ points, while yielding comparable results on pair classification and summarization tasks. This indicates that our method achieves stronger overall performance on unseen non-retrieval tasks and generates universal text embeddings without relying on task-related demonstrations. More importantly, bge-en-icl relies on knowledge distillation from a reranker model liu2025matryoshka as the teacher—an effective strategy that is not adopted by other baselines. In addition, incorporating in-context examples substantially increases the computational burden of LLMs, especially given that the maximum sequence length of bge-en-icl can reach up to 2048 tokens, which is four times that of our method, significantly limiting its applicability in resource-constrained scenarios. Under the compute-matched setting, our model achieves an average score improvement of $1.43$ over bge-en-icl (zero-shot). Moreover, bge-en-icl is sensitive to the selection strategy and number of in-context examples, hindering its applicability to different LLMs, especially for language models that lack in-context learning capabilities. Notably, our method is orthogonal to bge-en-icl. When equipped with the same ICL strategy, Causal2Vec (w/ ICL) achieves a state-of-the-art average score of 66.85, outperforming bge-en-icl (few-shot) by 0.77 points.
Figure 2: Average per-sample inference time (in milliseconds) and required sequence length of various Mistral-7B-based methods on selected MTEB subsets using a single A100 80GB GPU, including SprintDuplicateQuestions (Sprint.), RedditClusteringP2P (Reddit.), BiorxivClusteringP2P (Biorxiv.), etc. For the asymmetric dataset NFCorpus, we report the results per query-passage pair. LLM (Uni/Bi) denotes the standard Mistral-7B with causal or bidirectional attention.
Efficiency. In Figure 2, we report the approximate average inference time and required sequence length of different models on selected MTEB subsets. Specifically, ECHO and bge-en-icl inevitably increase inference time due to their extended sequence lengths. In contrast, since the additional bidirectional encoder contains only 110M parameters and a single Contextual token adds no computational burden to the LLM, Causal2Vec achieves inference speeds comparable to the standard Mistral-7B with causal or bidirectional attention, while reducing the required sequence length by up to 85% (Sprint.: 34.0 vs. 269.0; Emotion.: 62.1 vs. 421.9) and inference time by up to 82% (Sprint.: 4.37 vs. 26.09; Emotion.: 6.50 vs. 37.50) compared to bge-en-icl (few-shot).
4.5 Ablation Studies
4.5.1 Effectiveness of Each Component
To evaluate the effectiveness of the proposed Contextual token and representation method, we conduct ablation studies on MTEB-MINI using two base models of different scales: S-LLaMA-1.3B and Mistral-7B. As shown in Table 2, incorporating the Contextual token into LLM’s causal attention mechanism yields average score improvements of $0.24$ and $0.37$ for S-LLaMA-1.3B and Mistral-7B, respectively. These results not only confirm the effectiveness of the Contextual token but also highlight its scalability and applicability across different LLMs. We attribute these performance improvements to the rich contextualized content encoded in the Contextual token, which allows preceding tokens in the sequence to access accurate sentence information even without attending to future tokens, thereby mitigating the inherent architectural limitation in causal attention.
By concatenating the last hidden states of Contextual and EOS tokens as the final vector representation, we observe consistent performance improvements across all seven tasks, with average score gains of $0.72$ and $0.56$ on S-LLaMA-1.3B and Mistral-7B compared to standard LLMs, respectively. These results further confirm the generalizability and robustness of our method in enhancing embedding quality. We attribute the effectiveness of the proposed representation method to the following reasons: (1) it effectively alleviates the recency bias, as the Contextual token is not influenced by tokens near the end of sequence; (2) it enables explicit supervision of the Contextual token during training; and (3) the concatenation of two context-aware tokens enriches the semantic information of the final text embedding. See Appendix C.2 for further discussion on different representation methods.
Table 2: Performance comparison of different components on MTEB-MINI (41 datasets) using two base models: S-LLaMA-1.3B and Mistral-7B. CtxToken indicates adding the Contextual token to LLM’s input, while Concat denotes the proposed representation method that concatenates the last hidden states of LLM’s Contextual and EOS tokens as the text embedding.
CtxToken Concat Retr. (7) Rerank. (3) Clust. (6) PairClass. (3) Class. (11) STS (10) Summ. (1) Avg (41)
nDCG@10 MAP V-Meas. AP Acc. Spear. Spear.
S-LLaMA-1.3B
46.19 62.99 36.32 85.68 69.91 83.70 30.54 64.00
46.46 62.93 36.32 85.48 70.32 84.13 30.83 64.24
47.18 64.48 36.61 86.18 71.14 83.76 31.45 64.72
Mistral-7B
51.07 68.26 39.06 88.29 74.24 84.86 29.81 67.23
51.15 68.32 40.04 88.14 75.00 84.92 29.82 67.60
51.77 68.51 39.99 88.43 74.72 85.38 30.57 67.79
Table 3: Average MTEB-MINI score (41 datasets) for Causal2Vec with and without LoRA applied to the bidirectional encoder (Bi-LoRA). Refer to Appendix C.3 for detailed results for each task category.
Method S-LLaMA-1.3B Mistral-7B
w/ Bi-LoRA 64.72 67.66
w/o Bi-LoRA 64.43 67.79
Table 4: Average MTEB-MINI score (41 datasets) for placing the Contextual token before and after the task-specific instruction in Causal2Vec. See Appendix C.4 for detailed results for each task category.
Method S-LLaMA-1.3B Mistral-7B
before instruction 64.63 67.63
after instruction 64.72 67.79
4.5.2 Impact of Freezing the Bidirectional Encoder
Since the BERT-style bidirectional encoder (E5-base-v2) we use is specifically trained for embedding tasks, this section investigates whether it should be frozen during finetuning. As shown in Table 4, we observe that finetuning the bidirectional encoder with LoRA leads to an average score improvement of $0.29$ on the MTEB-MINI for S-LLaMA-1.3B, but degrades Mistral-7B’s performance by $0.13$ points. We attribute this to two potential effects of making the bidirectional encoder trainable: (1) it may cause catastrophic forgetting in the bidirectional encoder, and (2) it may help the LLM better interpret the Contextual token through joint finetuning. Large-scale LLMs are more susceptible to the former, as they already have sufficient capacity to comprehend newly added tokens during finetuning (llava, ; blip, ). This suggests that whether the introduced bidirectional encoder should remain frozen may depend on the scale of the underlying LLM.
4.5.3 The Position of Contextual Token
We investigate whether the Contextual token’s position affects embedding performance by comparing two placement settings: before vs. after the instruction. As shown in Table 4, "after instruction" consistently yields better results. We speculate that positioning the Contextual token before instruction tokens may hinder the LLM’s ability to accurately interpret and follow task-specific prompts.
4.5.4 The Number of Contextual Tokens
To examine the impact of using multiple Contextual tokens, we adopt cross-attention with a set of learnable queries to extract a fixed number of Contextual tokens from the bidirectional encoder, following detr ; blip2 . As presented in Table 5, increasing the number of Contextual tokens leads to performance degradation. We hypothesize that the additional tokens fail to provide more distinctive semantic information and instead introduce redundancy. These findings suggest that a single Contextual token is sufficient to supply the missing contextual information under causal attention, while maintaining model simplicity and efficiency.
Table 5: Performance comparison of Causal2Vec-S-LLaMA-1.3B using different numbers of Contextual tokens on the full MTEB (56 datasets). CtxToken denotes the Contextual token. Note: Causal2Vec uses a single Contextual token by default.
Task (# of datasets) Retr. (15) Rerank. (4) Clust. (11) PairClass. (3) Class. (12) STS (10) Summ. (1) Avg (56)
Metric nDCG@10 MAP V-Meas. AP Acc. Spear. Spear.
Causal2Vec 52.69 56.54 44.35 86.18 72.94 83.76 31.45 62.63
w/ 2 CtxTokens 52.48 56.27 44.37 86.06 72.43 83.90 31.60 62.47
w/ 4 CtxTokens 52.13 56.32 44.46 86.04 72.68 83.74 31.73 62.42
w/ 8 CtxTokens 52.46 56.28 44.44 85.99 72.56 83.75 30.67 62.46
4.5.5 Impact of Different Bidirectional Encoders
Table 6: Performance comparison of Causal2Vec-S-LLaMA-1.3B using different bidirectional encoders (Bi-Encoders) on the full MTEB (56 datasets). Note: E5-base-v2 is used as the default bidirectional encoder in Causal2Vec.
Task (# of datasets) Retr. (15) Rerank. (4) Clust. (11) PairClass. (3) Class. (12) STS (10) Summ. (1) Avg (56)
Metric nDCG@10 MAP V-Meas. AP Acc. Spear. Spear.
BERT-base bert 10.59 43.44 30.12 61.66 60.78 54.36 29.82 38.33
E5-small-v2 e5 49.04 54.32 39.92 84.67 72.94 80.39 31.16 59.93
E5-base-v2 e5 50.30 55.91 44.10 85.63 71.15 81.03 30.28 60.97
Causal2Vec 52.69 56.54 44.35 86.18 72.94 83.76 31.45 62.63
w/ E5-small-v2 52.45 55.82 44.15 86.08 72.65 83.92 31.24 62.43
w/ BERT-base 52.02 55.73 43.60 85.87 71.70 83.86 30.99 61.97
w/o Bi-Encoder 51.75 55.53 43.52 85.68 71.31 83.70 30.54 61.74
Finally, we explore the impact of different bidirectional encoders on MTEB-MINI. We conduct experiments using S-LLaMA-1.3B, with all bidirectional encoders being trainable to facilitate LLM’s adaptation to the Contextual token through joint finetuning. As shown in Table 6, incorporating even the standard BERT (bert, ) enhances embedding performance over the baseline LLM using last-token pooling (w/o Bi-Encoder). Moreover, the better-performing bidirectional encoder (e.g., E5-base vs. E5-small) generates a Contextual token with richer semantic information, enabling LLMs to capture more contextual semantics and thereby improve text embedding quality. It is important to note that although the E5-series encoders are specifically trained for embedding tasks, their performance on MTEB remains limited. This indicates that the performance improvements achieved by our method primarily stem from addressing the inherent shortcomings of causal attention and last-token pooling.
5 Conclusion
This paper presents Causal2Vec, a simple yet powerful text embedding model built upon decoder-only LLMs. It requires no architectural modifications or additional input text, achieving consistently strong performance in general-purpose embedding tasks. By introducing the proposed Contextual token, we enable each token in the sequence to capture contextual information within the inherent autoregressive modeling paradigm. To address the limitations of last-token pooling commonly used in unidirectional models, we propose a specialized representation method that concatenates the last hidden states of Contextual and EOS tokens as the final text embedding. Experimental results demonstrate that Causal2Vec not only achieves state-of-the-art performance on the MTEB benchmark, but also significantly reduces the required sequence length and inference time compared to best-performing methods.
Appendix A Experimental Details for Training
A.1 Hyperparameters
In this section, we provide additional training details based on Section 4.3. Specifically, we use the AdamW optimizer with an initial learning rate of 1e-4, and a warm-up strategy for the first 300 steps followed by linear decay over the remaining steps. To reduce GPU memory usage, all models are trained with bfloat16 quantization, gradient checkpointing, and FlashAttention-2 [54], following [11]. We also apply gradient accumulation to process a large batch size of 512 while sampling the same dataset within each batch. The maximum sequence length is set to the standard 512 tokens by default.
Table 7 presents the hyperparameters that vary across different base models. We set the LoRA rank to 64 and LoRA alpha to 32 for 7B models following [17], while using a large LoRA alpha of 128 for the 1.3B model. For the linear scheduler, the maximum steps are set to twice the training steps. Moreover, since we adopt the instruction-tuned versions of LLaMA-2-7B and Mistral-7B, the special [INST] and [/INST] tokens are appended to the input sequence following the official instruction prompt template.
A.2 Base Models
All base models and bidirectional encoders employed in this work are obtained from the Hugging Face platform, including: princeton-nlp/Sheared-LLaMA-1.3B, meta-llama/Llama-2-7b-chat-hf, mistralai/Mistral-7B-Instruct-v0.2, google-bert/bert-base-multilingual-cased, intfloat/e5-small-v2, and intfloat/e5-base-v2.
A.3 Public Retrieval Datasets and Instructions
The collection of publicly available retrieval datasets used for training is curated by [16] and includes the following datasets: ELI5 (sample ratio 0.1) [55], HotpotQA [56], FEVER [57], MIRACL [58], MS-MARCO passage ranking (sample ratio 0.5) and document ranking (sample ratio 0.2) [59], NQ [3], NLI [21], SQuAD [60], TriviaQA [61], Quora Duplicate Questions (sample ratio 0.1) [62], Mr. TyDi [63], DuReader [64], and T2Ranking (sample ratio 0.5) [65].
Table 8 lists the instructions used for each dataset, which are manually written by [32]. Notably, in the query-passage pairs of retrieval datasets, task-specific instructions are appended only to the queries, without modifying the passages.
Table 7: Hyperparameters used in the experiments.
Hyperparameter S-LLaMA-1.3B Mistral-7B & LLaMA-2-7B
Batch Size 128 64
Gradient Accumulation Steps 4 8
Training Steps 2000 1000
Maximum Steps 4000 2000
LoRA Rank 64 64
LoRA Alpha 128 32
LoRA for Bidirectional Encoder
Table 8: Instructions used for public retrieval datasets.
Dataset Instruction (s)
ELI5 Provided a user question, retrieve the highest voted answers on Reddit ELI5 forum
HotpotQA Given a multi-hop question, retrieve documents that can help answer the question
FEVER Given a claim, retrieve documents that support or refute the claim
MIRACL Given a question, retrieve Wikipedia passages that answer the question
MSMARCO Passage Given a web search query, retrieve relevant passages that answer the query
MSMARCO Document Given a web search query, retrieve relevant documents that answer the query
NQ Given a question, retrieve Wikipedia passages that answer the question
NLI Given a premise, retrieve a hypothesis that is entailed by the premise
Retrieve semantically similar text
SQuAD Retrieve Wikipedia passages that answer the question
TriviaQA Retrieve Wikipedia passages that answer the question
QuoraDuplicates Given a question, retrieve questions that are semantically equivalent to the given question
Find questions that have the same meaning as the input question
Mr. TyDi Given a question, retrieve Wikipedia passages that answer the question
DuReader Given a Chinese search query, retrieve web passages that answer the question
T2Ranking Given a Chinese search query, retrieve web passages that answer the question
Appendix B Experimental Details for Evaluation
B.1 MTEB-MINI Details
Considering the substantial computational resources required for full evaluation on MTEB, we follow [11, 16] and select a subset of the MTEB for ablation and analysis. While prior studies utilize only a few datasets, our preliminary experiments suggest that evaluation on a limited subset may introduce significant bias and fail to effectively reflect the overall trends of the full MTEB. We empirically argue that a representative subset should cover as many MTEB datasets as possible to ensure consistency with evaluation results of the complete MTEB. To this end, as shown in Table 9, we introduce the MTEB-MINI by selecting 41 datasets spanning all task categories in MTEB.
Table 9: Composition of the MTEB-MINI benchmark.
Category Dataset
Retrieval (7) ArguAna, SciFact, NFCorpus, FiQA2018, SCIDOCS, TRECCOVID, Touche2020
Reranking (3) AskUbuntuDupQuestions, SciDocsRR, StackOverflowDupQuestions
Clustering (6) BiorxivClusteringS2S, BiorxivClusteringP2P, MedrxivClusteringS2S, MedrxivClusteringP2P
MedrxivClusteringP2P, TwentyNewsgroupsClustering, StackExchangeClusteringP2P
Pair Classification (3) SprintDuplicateQuestions, TwitterSemEval2015, TwitterURLCorpus
Classification (11) AmazonCounterfactualClassification, AmazonReviewsClassification, Banking77Classification
EmotionClassification, MassiveIntentClassification, MassiveScenarioClassification
MTOPDomainClassification, MTOPIntentClassification, ToxicConversationsClassification
TweetSentimentExtractionClassification, ImdbClassification
STS (10) BIOSSES, SICK-R, STS12, STS13, STS14, STS15, STS16, STS17, STS22, STSBenchmark
SummEval (1) SummEval
Overall 41 datasets
B.2 Instructions for MTEB Evaluation
To enable a fair comparison with prior leading embedding methods [32, 16, 11, 13, 17], we use the same instruction prompts for evaluation on both MTEB and MTEB-MINI. The instructions applied to each dataset are listed in Table 13.
Appendix C Additional Results
C.1 The L2 Norms of Contextual and EOS Tokens
To examine the respective contributions of Contextual and EOS tokens to the final text embedding, we compare their L2 norms on selected MTEB datasets. As depicted in Figure 3, we observe that the EOS token consistently shows higher L2 norms across various task categories, indicating its greater influence on the concatenated representation.
(a) S-LLaMA-1.3B
(b) Mistral-7B
Figure 3: L2 norms of Contextual and EOS tokens on selected MTEB subsets for two base models: S-LLaMA-1.3B and Mistral-7B. The evaluated datasets span seven tasks, including NFCorpus, SciDocsRR, BiorxivClusteringP2P (Biorxiv.), SprintDuplicateQuestions (Sprint.), EmotionClassification (Emotion.), STS12, and SummEval.
C.2 Different Representation Methods
We also explore different representation methods tailored to our embedding framework. As shown in Table 10, both concatenation strategies consistently outperform last-token pooling for S-LLaMA-1.3B and Mistral-7B. This suggests that incorporating an additional context-aware token with the EOS token leads to richer semantic information while reducing the model’s reliance on the single EOS token alone. Additionally, we observe further performance improvements when the concatenated Contextual token is derived from the LLM, rather than from the bidirectional encoder followed by an MLP layer. We speculate that this helps the LLM better capture the semantic content encoded in the Contextual token.
Table 10: Performance comparison on MTEB-MINI (41 datasets) using different representation method with two base models: S-LLaMA-1.3B and Mistral-7B. By default, Causal2Vec generates text embedding by concatenating the output hidden states of LLM’s Contextual and EOS tokens, while Bi-EOS denotes concatenating LLM’s output EOS token with the bidirectional encoder’s output that has been processed by an MLP layer. Note: all experiments incorporate the Contextual token into LLM’s input sequence.
Task (# of datasets) Retr. (7) Rerank. (3) Clust. (6) PairClass. (3) Class. (11) STS (10) Summ. (1) Avg (41)
Metric nDCG@10 MAP V-Meas. AP Acc. Spear. Spear.
S-LLaMA-1.3B
Causal2Vec 47.18 64.48 36.61 86.18 71.14 83.76 31.45 64.72
w/ Bi-EOS 46.57 62.92 36.70 85.19 70.76 84.05 30.44 64.39
w/ last-token 46.46 62.93 36.32 85.48 70.32 84.13 30.83 64.24
Mistral-7B
Causal2Vec 51.77 68.51 39.99 88.43 74.72 85.38 30.57 67.79
w/ Bi-EOS 51.50 68.38 39.73 88.37 74.90 85.17 30.03 67.68
w/ last-token 51.15 68.32 40.04 88.14 75.00 84.92 29.82 67.60
C.3 Impact of Freezing the Bidirectional Encoder
To investigate the impact of freezing the bidirectional encoder during Causal2Vec finetuning, we compare the performance of applying "w/ Bi-LoRA" and "w/o Bi-LoRA" with two base models: S-LLaMA-1.3B, and Mistral-7B. Table 11 presents the detailed results across all seven tasks on MTEB-MINI.
Table 11: Performance comparison on MTEB-MINI (41 datasets) with and without applying LoRA to the bidirectional encoder (Bi-LoRA), using two base models: S-LLaMA-1.3B and Mistral-7B.
Task (# of datasets) Retr. (7) Rerank. (3) Clust. (6) PairClass. (3) Class. (11) STS (10) Summ. (1) Avg (41)
Metric nDCG@10 MAP V-Meas. AP Acc. Spear. Spear.
S-LLaMA-1.3B
w/ Bi-LoRA 47.18 64.48 36.61 86.18 71.14 83.76 31.45 64.72
w/o Bi-LoRA 46.80 63.97 36.26 86.15 70.57 83.92 30.76 64.43
Mistral-7B
w/ Bi-LoRA 51.50 68.23 39.64 88.43 74.86 85.20 30.41 67.66
w/o Bi-LoRA 51.77 68.51 39.99 88.43 74.72 85.38 30.57 67.79
Table 12: Performance comparison between placing the Contextual token before and after instruction on MTEB-MINI (41 datasets) using two base models: S-LLaMA-1.3B and Mistral-7B.
Task (# of datasets) Retr. (7) Rerank. (3) Clust. (6) PairClass. (3) Class. (11) STS (10) Summ. (1) Avg (41)
Metric nDCG@10 MAP V-Meas. AP Acc. Spear. Spear.
S-LLaMA-1.3B
before instruction 47.17 64.18 36.78 86.10 70.89 83.76 30.70 64.63
after instruction 47.18 64.48 36.61 86.18 71.14 83.76 31.45 64.72
Mistral-7B
before instruction 51.65 68.16 39.26 88.21 74.96 85.23 29.55 67.63
after instruction 51.77 68.51 39.99 88.43 74.72 85.38 30.57 67.79
C.4 The Position of Contextual Token
In Table 12, we report the detailed MTEB-MINI results for different positions of the Contextual token, including before and after the instruction tokens.
C.5 Full MTEB Results
We present detailed results on all 56 MTEB datasets for the proposed Causal2Vec in Table 14, including three base models: S-LLaMA-1.3B, LLaMA-2-7B, and Mistral-7B.
Appendix D Limitations
Despite the effectiveness of Causal2Vec, several limitations should be acknowledged: (1) Our findings suggest that a single Contextual token is sufficient to provide the missing contextual information for decoder-only LLMs. Future work could explore generating additional Contextual tokens using different bidirectional encoders or utilizing multiple task-related examples. (2) Our experiments are limited to three popular LLMs with fewer than 7B parameters, while further validation on more diverse and larger-scale LLMs could better demonstrate the scalability and robustness of our proposed mechanism.
Appendix E Broader Impacts
Our proposed Causal2Vec can be applied to a wide range of real-world applications, including information retrieval and LLM-based retrieval-augmented generation systems. However, LLMs suffer from biases and hallucinations, which could potentially lead to negative societal impacts.
Table 13: Instructions used for evaluation on the MTEB. “STS*” denotes that the corresponding instruction is applied to all STS datasets.
Dataset Instruction Template
AmazonCounterfactualClassification
Classify a given Amazon customer review text as either counterfactual
or not-counterfactual.
AmazonPolarityClassification Classify Amazon reviews into positive or negative sentiment
AmazonReviewsClassification Classify the given Amazon review into its appropriate rating category
Banking77Classification Given a online banking query, find the corresponding intents
EmotionClassification
Classify the emotion expressed in the given Twitter message into one of the six
emotions: anger, fear, joy, love, sadness, and surprise.
ImdbClassification Classify the sentiment expressed in the given movie review text from the IMDB dataset.
MassiveIntentClassification Given a user utterance as query, find the user intents
MassiveScenarioClassification Given a user utterance as query, find the user scenarios
MTOPDomainClassification Classify the intent domain of the given utterance in task-oriented conversation
MTOPIntentClassification Classify the intent of the given utterance in task-oriented conversation
ToxicConversationsClassif. Classify the given comments as either toxic or not toxic
TweetSentimentClassification Classify the sentiment of a given tweet as either positive, negative, or neutral
ArxivClusteringP2P Identify the main and secondary category of Arxiv papers based on the titles and abstracts.
ArxivClusteringS2S Identify the main and secondary category of Arxiv papers based on the titles
BiorxivClusteringP2P Identify the main category of Biorxiv papers based on the titles and abstracts
BiorxivClusteringS2S Identify the main category of Biorxiv papers based on the titles
MedrxivClusteringP2P Identify the main category of Medrxiv papers based on the titles and abstracts
MedrxivClusteringS2S Identify the main category of Medrxiv papers based on the titles
RedditClustering Identify the topic or theme of Reddit posts based on the titles
RedditClusteringP2P Identify the topic or theme of Reddit posts based on the titles and posts
StackExchangeClustering Identify the topic or theme of StackExchange posts based on the titles
StackExchangeClusteringP2P Identify the topic or theme of StackExchange posts based on the given paragraphs
TwentyNewsgroupsClustering Identify the topic or theme of the given news articles
SprintDuplicateQuestions Retrieve duplicate questions from Sprint forum
TwitterSemEval2015 Retrieve tweets that are semantically similar to the given tweet
TwitterURLCorpus Retrieve tweets that are semantically similar to the given tweet
AskUbuntuDupQuestions Retrieve duplicate questions from AskUbuntu forum
MindSmallReranking Retrieve relevant news articles based on user browsing history
SciDocsRR Given a title of a scientific paper, retrieve the titles of other relevant papers
StackOverflowDupQuestions Retrieve duplicate questions from StackOverflow forum
ArguAna Given a claim, find documents that refute the claim
ClimateFEVER Given a claim about climate change, retrieve documents that support or refute the claim.
CQADupstackRetrieval
Given a question, retrieve detailed question descriptions from Stackexchange that are
duplicates to the given question.
DBPedia Given a query, retrieve relevant entity descriptions from DBPedia
FEVER Given a claim, retrieve documents that support or refute the claim
FiQA2018 Given a financial question, retrieve user replies that best answer the question
HotpotQA Given a multi-hop question, retrieve documents that can help answer the question
MSMARCO Given a web search query, retrieve relevant passages that answer the query
NFCorpus Given a question, retrieve relevant documents that best answer the question
NQ Given a question, retrieve Wikipedia passages that answer the question
QuoraRetrieval Given a question, retrieve questions that are semantically equivalent to the given question.
SCIDOCS Given a scientific paper title, retrieve paper abstracts that are cited by the given paper
SciFact Given a scientific claim, retrieve documents that support or refute the claim
Touche2020 Given a question, retrieve detailed and persuasive arguments that answer the question
TRECCOVID Given a query on COVID-19, retrieve documents that answer the query
STS* Retrieve semantically similar text.
BUCC/Tatoeba Retrieve parallel sentences.
SummEval Given a news summary, retrieve other semantically similar summaries
Table 14: Results of Causal2Vec on all 56 MTEB datasets across three base models: S-LLaMA-1.3B, LLaMA-2-7B, and Mistral-7B.
Dataset S-LLaMA-1.3B LLaMA-2-7B Mistral-7B
Mistral-7B
(w/ ICL)
AmazonCounterfactualClassification 74.49 76.79 76.22 75.99
AmazonPolarityClassification 92.75 94.80 95.02 95.80
AmazonReviewsClassification 46.48 51.75 51.40 53.78
ArguAna 54.73 57.35 57.55 59.11
ArxivClusteringP2P 46.25 48.37 48.99 50.64
ArxivClusteringS2S 39.52 42.88 45.51 47.09
AskUbuntuDupQuestions 61.59 63.54 65.96 65.71
BIOSSES 83.96 84.06 86.42 87.28
Banking77Classification 85.96 88.14 88.62 88.85
BiorxivClusteringP2P 38.13 39.05 39.24 40.82
BiorxivClusteringS2S 35.13 36.42 38.32 39.09
CQADupstackRetrieval 39.53 43.42 45.59 46.82
ClimateFEVER 32.62 32.46 35.55 34.12
DBPedia 44.85 49.85 51.65 52.09
EmotionClassification 46.82 49.74 50.56 51.40
FEVER 88.11 90.53 91.53 91.68
FiQA2018 44.52 51.29 54.96 56.03
HotpotQA 67.13 71.45 74.25 73.11
ImdbClassification 83.76 88.33 91.24 92.45
MSMARCO 40.88 41.22 42.22 42.83
MTOPDomainClassification 94.05 95.53 95.79 96.36
MTOPIntentClassification 73.45 82.38 83.12 86.23
MassiveIntentClassification 73.36 77.63 78.14 79.53
MassiveScenarioClassification 77.58 79.88 81.27 82.40
MedrxivClusteringP2P 33.38 33.13 34.33 36.32
MedrxivClusteringS2S 31.58 32.14 34.28 34.73
MindSmallReranking 32.71 32.46 32.32 32.31
NFCorpus 37.43 40.21 41.63 41.41
NQ 58.02 64.10 66.65 66.50
QuoraRetrieval 88.91 88.80 89.35 89.24
RedditClustering 56.91 63.07 64.73 64.99
RedditClusteringP2P 61.26 64.31 66.43 68.21
SCIDOCS 19.56 21.28 22.40 22.76
SICK-R 81.99 82.78 83.49 83.33
STS12 77.04 78.77 79.37 79.71
STS13 87.47 88.89 88.69 89.75
STS14 83.21 85.29 85.43 85.80
STS15 88.93 89.86 90.76 90.92
STS16 86.83 87.72 88.26 88.38
STS17 91.13 92.19 92.47 92.31
STS22 69.21 70.67 69.44 69.20
STSBenchmark 87.85 88.77 89.47 89.96
SciDocsRR 81.68 84.11 84.40 84.61
SciFact 73.04 75.77 77.52 77.92
SprintDuplicateQuestions 96.26 97.00 96.70 97.16
StackExchangeClustering 64.27 69.14 72.23 76.40
StackExchangeClusteringP2P 32.41 36.74 37.73 40.63
StackOverflowDupQuestions 50.16 52.61 55.16 54.81
SummEval 31.45 31.09 30.57 30.82
TRECCOVID 76.24 78.65 83.48 83.09
Touche2020 24.74 22.83 24.86 25.51
ToxicConversationsClassification 65.03 65.02 63.05 65.14
TweetSentimentExtractionClassification 61.53 61.44 62.52 62.46
TwentyNewsgroupsClustering 49.01 54.33 56.01 59.62
TwitterSemEval2015 75.24 79.78 81.35 82.91
TwitterURLCorpus 87.03 86.77 87.23 87.50
MTEB Average (56) 62.63 64.94 66.10 66.85